OpenStudy (anonymous):

Cool treats sold 60 ice cream cones.Single dip cones sold for $2.50 each and double dip cones for $4.15 each.In all $179.70 was taken in for the cones.How many of each size cone were sold?.I was helped for most of it but am stuck at the multiplying part.

OpenStudy (dan815):

lets s be for the number of single dip cones sold let d be the number of double dip cones sold The information give is that 60 total cones were sold s+d=60 price for each single = 2.50 price for each double = 4.15 total price = 179.70 2.50*s + 4.15*d=179.70 ------------------------------------- You have these 2 equations s+d=60 2.50*s + 4.15*d=179.70 solve for s and d

OpenStudy (dan815):

does it make senseÉ

OpenStudy (anonymous):

somewhat.. I am following.

OpenStudy (dan815):

To clear things up any further, the price for s number of single dip cones would be 2.50 * s as you pay 2.50 for each single dip so, if there are s of those, single dips then you pay 2.50 s times similiary price for all the double dip cones together is 4.15*d

OpenStudy (dan815):

so we get this equation representing the total sales 2.50*s+4.15*d=179.70

OpenStudy (anonymous):

ok I get this part ..How do I go about solving the equation